OpenStudy (anonymous):

What is the slope intercept of -6x+y=9?

OpenStudy (anonymous):

m=9 So the slope intercept is k=6

OpenStudy (anonymous):

So is it y=6x+9? Or y=6x-9?

OpenStudy (anonymous):

The first: y=6x+9

OpenStudy (anonymous):

Thankkk youu(:

OpenStudy (amistre64):

somethings wrong with that ...

OpenStudy (amistre64):

slope intercept form is: y = mx + b , such that m is the slope and b is the intercept; compare that to the answer provided and see if you can notice the error

OpenStudy (anonymous):

In sweden we have y=kx+m

OpenStudy (amistre64):

sweden eh ... then that makes sense ;)